Font definition option cmmib57.sty

Michael Downes

Version 2.2, 1995/01/26

1 Introduction

This file contains alternate font definition statements for the fonts cmmib and cmbsy that are tailored to the set of sizes provided in the Y&Y/Blue Sky Research PostScript versions of the AMSFonts collection. In that font set, the only .tfm files provided below size 10 are 5 and 7, rather than 5,6,7,8,9. This means that if the standard .fd files are used, any attempt to use 6,8,9 point math will result in error messages about missing .tfm files.

Rather than provide substitute versions of the standard OMLcmm.fd and OMScmsy.fd font definition files, we put the alternate font definitions in package form so that any document that uses them will have an explicit indication in the document:

 \usepackage{cmmib57}

Scaling from sizes 5 and 7 to the surrounding sizes results in some font metric differences, so the line breaks (and, consequently, page breaks) resulting from that alternative won’t necessarily be the same as with the standard L^{A}TEX font definition files.

Package info:

  \ProvidesPackage{cmmib57.sty}[1995/01/26 v2.2 %
    alternate cmmib, cmbsy font definitions]

We assume \DeclareFontFamily statements have already been issued for the cmm and cmsy families.

  \DeclareFontShape{OML}{cmm}{b}{n}{
    <-6> cmmib5 <6-8> cmmib7 <8-> cmmib10
  }{}
  \DeclareFontShape{OMS}{cmsy}{b}{n}{
    <-6> cmbsy5 <6-8> cmbsy7 <8-> cmbsy10
  }{}